2. Graph the isoclines of the ODE Y = y + t2 where the slope is -1,0,1.
Answer #1

Clower Given thati- where the slope is -1,0,1.. yfcast) then i soclines m=y= f(xt) at slope m will be fra,t)=m we have slopegraph will be NOW, at m=o, g+t² m ytte: 0 y=-62 Similarly for m at 1: Ytt? - m Ytt? - 1 y = 1-12 y=1-423